# Controllable Video Generation with Provable Disentanglement

### v2

@@ B.1 Controllable Video Generation
Recent advances in controllable video generation have led to significant progress, with text-to-video (T2V) models [51, 34, 8, 57, 56] achieving impressive results in generating videos from textual descriptions. However, effectiveness of the control is highly dependent on the quality of the input prompt, making it difficult to achieve fine-grained control over the generated content. An alternative method for control involves leveraging side information such as pose [36, 59], camera motion [50, 45], depth [27, 48] and so on. While this approach allows for more precise control, it requires paired data, which can be challenging to collect. Besides, most of the aforementioned alignment-based techniques share a common issue: the control signals are directly aligned with the entire video. This issue not only reduces efficiency but also complicates the task of achieving independent control over different aspects of the video, which further motivates us to propose a framework to find the disentanglement representation for conditional generation.
@@ B.2 Nonlinear Independent Component Analysis
Nonlinear independent component analysis offers a potential approach to uncover latent causal variables in time series data. These methods typically utilize auxiliary information, such as class labels or domain-specific indices, and impose independence constraints to enhance the identifiability of latent variables. Time-contrastive learning (TCL) [12] builds on the assumption of independent sources and takes advantage of the variability in variance across different segments of data. Similar Permutation-based contrastive learning (PCL) [13] introduces a learning framework that tell true independent sources from their permuted counterparts. Additionally, i-VAE [16] employs deep neural networks and Variational Autoencoders (VAEs) to closely approximate the joint distribution of observed data and auxiliary non-stationary regimes. Recently, [53, 52] extends the identifiability to linear and nonlinear non-Gaussian cases without auxiliary variables, respectively. CaRiNG [2] further tackles the case when the mixing process is non-invertible. Additionally, CITRIS [26, 25] emphasizes the use of intervention target data, and IDOL [22] incorporates sparsity into the latent transition process to identify latent variables, even in the presence of instantaneous effects.
Appendix C Datasets Details
@@
FaceForensics [32] is a forensics dataset consisting of 1000 original video sequences all videos contain a trackable mostly frontal face without occlusions. SkyTimelapse [49] typically consists of sequential images or videos capturing the dynamic behavior of the sky over time.
@@
RealEstate10K [58] is a large dataset of camera poses corresponding to 10 million frames derived from about 80,000 video clips, gathered from about 10,000 YouTube videos. For each clip, the poses form a trajectory where each pose specifies the camera position and orientation along the trajectory. These poses are derived by running SLAM and bundle adjustment algorithms on a large set of videos. To the best of our knowledge, the proposed method is the first GAN-based approach to leverage the complex camera pose dataset for validating unconditional video generation.
Appendix D Reproductability
@@
All of our models are trained on an NVIDIA A100 40G GPU. Our model requires around a throughput of 20 million frames for convergence. For the baseline models, we use the official implementation with the default hyperparameters. The configuration of hyperparameters for our model training is as shown in Table  4.
@@
Dimensionality of ${\mathbf{z}}_{s}^{t}$ FaceForensics: 4, SkyTimelapse: 12, RealEstate: 12
@@
Table 5: Architecture of Temporal Transition Module
@@
Appendix E More Experiments
E.1 Computational Efficiency.
@@ E.2 More Videos
Appendix F Impact Statements
@@
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: A detailed list of contribution can be found at the end of Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Assumptions mentioned in the theorem part are discussed in the ’Illustration of assumption’ part and Appendix A.2. A discussion of limitation can be found at the end of Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Proof of the theorem can be found in Appendix A.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The code is provided in the supplemental material. The details of the implementation can be found in the Appendix D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We release the code in the supplemental materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting can be found in the Section 5. The details of the implementation can be found in the Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive for a video generative model.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The required computing resources are provided in the Appendix D and E.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We follow the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The impact statement can be found in Appendix F.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original paper of baselines and datasets used in this paper are properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v3

@@ B.1 Controllable Video Generation
Recent advances in controllable video generation have led to significant progress, with text-to-video (T2V) models (Yang et al., 2024b; Singer et al., 2022; Ho et al., 2022; Zhou et al., 2022; Zheng et al., 2024) achieving impressive results in generating videos from textual descriptions. However, effectiveness of the control is highly dependent on the quality of the input prompt, making it difficult to achieve fine-grained control over the generated content. An alternative method for control involves leveraging side information such as pose (Tu et al., 2024; Zhu et al., 2025), camera motion (Yang et al., 2024a; Wang et al., 2024), depth (Liu et al., 2024; Xing et al., 2024) and so on. While this approach allows for more precise control, it requires paired data, which can be challenging to collect. Besides, most of the aforementioned alignment-based techniques share a common issue: the control signals are directly aligned with the entire video. This issue not only reduces efficiency but also complicates the task of achieving independent control over different aspects of the video, which further motivates us to propose a framework to find the disentanglement representation for conditional generation.
@@ B.2 Nonlinear Independent Component Analysis
Nonlinear independent component analysis offers a potential approach to uncover latent causal variables in time series data. These methods typically utilize auxiliary information, such as class labels or domain-specific indices, and impose independence constraints to enhance the identifiability of latent variables. Time-contrastive learning (TCL) (Hyvarinen & Morioka, 2016) builds on the assumption of independent sources and takes advantage of the variability in variance across different segments of data. Similar Permutation-based contrastive learning (PCL) (Hyvarinen & Morioka, 2017) introduces a learning framework that tell true independent sources from their permuted counterparts. Additionally, i-VAE (Khemakhem et al., 2020) employs deep neural networks and Variational Autoencoders (VAEs) to closely approximate the joint distribution of observed data and auxiliary non-stationary regimes. Recently, (Yao et al., 2021; 2022) extends the identifiability to linear and nonlinear non-Gaussian cases without auxiliary variables, respectively. CaRiNG (Chen et al., 2024) further tackles the case when the mixing process is non-invertible. Additionally, CITRIS (Lippe et al., 2022; 2023) emphasizes the use of intervention target data, and IDOL (Li et al., 2024c) incorporates sparsity into the latent transition process to identify latent variables, even in the presence of instantaneous effects.
Appendix C Dataset Details
We use the following datasets to verify our model.
@@ Appendix C Dataset Details
  • FaceForensics (Rössler et al., 2018): A forensics dataset consisting of 1,000 original video sequences. All videos contain a trackable, mostly frontal face without occlusions.
  • SkyTimelapse (Xiong et al., 2018): Typically consists of sequential images or videos capturing the dynamic behavior of the sky over time. We use the 2,368 officially released videos.
@@
  • RealEstate10K (Zhou et al., 2018): A large dataset of camera poses corresponding to 10 million frames from about 80,000 video clips, gathered from around 10,000 YouTube videos. Each clip’s poses form a trajectory specifying camera position and orientation, derived via SLAM and bundle adjustment. To our knowledge, our method is the first GAN-based approach to leverage this dataset for unconditional video generation.
  • CelebV-HQ (Zhu et al., 2022): A large-scale, high-quality video dataset with diverse celebrity identities and actions. It contains 35,666 video clips with a minimum resolution of 512×512, covering 15,653 identities.
Appendix D Metric Details
We use the following metrics to verify our model.
  • FVD (Fréchet Video Distance (Unterthiner et al., 2019)) is a metric for evaluating video generation quality. It compares real and generated videos by extracting spatiotemporal features with a pretrained model and computing the Fréchet distance between their feature distributions.
  • SAP (Separated Attribute Predictability (Kumar et al., 2017)) measures how well each latent dimension is associated with a single ground-truth factor. It’s computed by training a simple regressor to predict true factors from each latent dimension, then comparing the two best-performing dimensions. A higher score means better disentanglement.
  • Modularity (Ridgeway & Mozer, 2018) measures whether each latent dimension encodes information about at most one true factor. It penalizes when a single latent variable carries mixed information about multiple true factors. Perfect modularity means each latent dimension corresponds only to a single factor.
  • MCC (Mean Correlation Coefficient) is a commonly used disentanglement metric in representation learning. Let $Z\in\mathbb{R}^{D}$ be the ground-truth latent vector and $\widehat{Z}\in\mathbb{R}^{\widehat{D}}$ the estimated vector. To calculate MCC, we first compute the Pearson correlations $R_{ij}=\mathrm{corr}(Z_{i},\widehat{Z}_{j}),$ then select an injective matching $\pi:\{1,\dots,D\}\to\{1,\dots,\widehat{D}\}$ maximizing $\sum_{i=1}^{D}|R_{i,\pi(i)}|$. Finally, the MCC value is defined as $\mathrm{MCC}=\frac{1}{D}\sum_{i=1}^{D}\bigl|R_{i,\pi(i)}\bigr|.$
Appendix E Reproductability
@@
All of our models are trained on an NVIDIA A100 40G GPU. Our model requires around a throughput of 20 million frames for convergence. For the baseline models, we use the official implementation with the default hyperparameters. The configuration of hyperparameters for our model training is as shown in Table 4.
@@
Dimensionality of ${\mathbf{z}}_{s}^{t}$ - FaceForensics: 4, SkyTimelapse: 12, RealEstate: 12
@@
Table 5: Architecture of Temporal Transition Module.
@@
Appendix F More Experiments
F.1 Longer Videos
We also verify our methods on generating longer videos with 32 frames on the FaceForensics Dataset, as shown in Table 6. Our method achieves substantially better performance than MoStGAN-V on disentanglement metrics.
Table 6: Performance on FaceForensics with length=32.
Metrics MoStGAN-V CoVoGAN
FVD $\downarrow$ 159.24 145.77
MCC (%) $\uparrow$ 14.23 29.99
SAP (%) $\uparrow$ 0.87 4.20
Modularity (%) $\uparrow$ 8.15 10.14
F.2 Choice of $n_{s}$
In CoVoGAN, the dimension of $n_{s}$ is treated as a tunable hyperparameter. While the choice of $n_{s}$ is important, it turns out to be not overly sensitive. From a theoretical standpoint, assuming a known latent dimensionality is standard in the nonlinear ICA literature (Hyvarinen & Morioka, 2016; Kong et al., 2023). Moreover, even when $n_{s}$ is not specified exactly, the model can still yield meaningful results: if $n_{s}$ is set too large, the model tends to fit some noise, whereas if it is set too small, certain components may be merged. This explains why our empirical results show robustness with respect to $n_{s}$. As presented in Table 7, varying it within a reasonable range does not lead to significant performance changes.
Table 7: Ablation on $n_{s}$ on the FaceForensics dataset.
$n_{s}$ 4 8
FVD $\downarrow$ 48.8 47.9
F.3 Supervised Video Generation
In this paper, we propose a method for controlling video generation without relying on specific supervisory signals such as text or trajectories. Our goal is not to surpass supervised approaches like text-to-video or trajectory-guided models, but rather to explore the feasibility of achieving disentangled video generation in the absence of explicit supervision. This setting is considerably more challenging, as annotations are typically task-specific and costly to obtain. Moreover, when appropriate supervisory signals such as text or trajectories are available, our method can be combined with them to enhance controllability through an identifiable generative process.
To further strengthen our evaluation, we have integrated our TTM module into two additional models: the diffusion-based Text-to-Video generation model, Wan (Wan et al., 2024), and the trajectory-supervised model, MotionCtrl (Wang et al., 2024).
For text-to-video generation, we integrate our TTM module into the Wan-1.3B model and finetune it on a subset of WebVid-10M (Bain et al., 2021), using the first 10% of the data. We conduct a thorough comparison against the original Wan-1.3B model and CogVideoX-2B (Yang et al., 2024b). The results, presented in Table 8, were evaluated using the VBench (Huang et al., 2023) toolkit for the first six metrics, and three different implementations of CLIP similarity: hf_clip_vit_b32, hf_clip_vit_l14 (Radford et al., 2021), hf_laion_clip_vit_b32 (Rössler et al., 2018) for the controllability measures. While our model shows a slight decrease in some metrics compared to CogVideoX-2B (e.g., hf_clip_vit_b32), it still outperforms the original Wan-1.3B model, especially to the aspect of controllability. Examples of generated videos are provided in Figure 9.
Table 8: Comparison of text-to-video generation across multiple metrics (larger is better).
Metric CogVideoX-2B Wan-1.3B TTM + Wan-1.3B
imaging_quality $\uparrow$ 0.4870 0.6851 0.6877
motion_smoothness $\uparrow$ 0.9886 0.9874 0.9935
dynamic_degree $\uparrow$ 0.3465 0.3960 0.4059
subject_consistency $\uparrow$ 0.9347 0.9324 0.9529
background_consistency $\uparrow$ 0.9585 0.9375 0.9628
aesthetic_quality $\uparrow$ 0.3849 0.5370 0.4924
hf_clip_vit_b32 $\uparrow$ 0.2428 0.2234 0.2288
hf_clip_vit_l14 $\uparrow$ 0.2007 0.1788 0.1892
hf_laion_clip_vit_b32 $\uparrow$ 0.1859 0.1791 0.1980
We also integrate our TTM module into the MotionCtrl model and conducted a comparison on the RealEstate10K dataset, with regard to video generation conditioned on trajectory of camera poses. As shown in Table 9, our TTM module successfully improves the performance of the MotionCtrl model.
Table 9: Comparison of MotionCtrl and TTM + MotionCtrl across multiple metrics (lower is better).
Metric MotionCtrl TTM + MotionCtrl
CamMC $\downarrow$ 0.0840 0.0776
FID $\downarrow$ 130.29 129.05
FVD $\downarrow$ 934.37 917.28
Since both baseline models are diffusion-based rather than GAN-based, we integrate the TTM module through cross-attention as a conditioning mechanism, following a similar strategy to that used in an identifiable image generation model (Xie et al., 2025). Although this design is not fully aligned with our theoretical framework, it highlights a promising direction for future investigation.
F.4 Computational Efficiency
@@ F.5 More Visualization Results
F.6 Experiments on Human Motion Dataset
We also conduct experiments on more TaiChi (Sun et al., 2017), a more complex dataset with richer human motion dynamics. We compare FVD, MCC, SAP, and Modularity on the TaiChi $64^{2}$ dataset. For the disentanglement metrics, we use the open-source package MediaPipe to extract human motion keypoints as ground-truth variables. However, for MoStGAN-V, MediaPipe fails to extract any keypoints from the generated videos due to poor generation quality, making the disentanglement metrics inapplicable for this method. The quality results and quantity results can be found in Figure 8 and Table 11 respectively.
Table 11: Comparison of StyleGAN-V, MoStGAN-V, and CoVoGAN on Taichi.
Metrics on Taichi StyleGAN-V MoStGAN-V CoVoGAN
FVD $\downarrow$ 68.52 60.39 55.3
MCC (%) $\uparrow$ 18.2 - 42.3
SAP (%) $\uparrow$ 1.5 - 3.2
Modularity (%) $\uparrow$ 0.2 - 1.9
F.7 Experiments on Independent Utility of $z^{c}$ and $z^{s}$
In this subsection, we verify that the $z^{c}$ and $z^{s}$ can be utilized independently to generate videos. We have added the following ablation experiments: we first randomly sample $z^{c}$ and $z^{s}$ independently from the learned distribution, then for each $z^{c}$, we use different $z^{s}$ to generate videos, and vice versa. We compare the FVD and disentanglement metrics and both show that the two representations can be used independently to generate videos with good performance. The results are shown in Table 12.
Table 12: Extrapolation ablation study on FaceForensics.
Metrics Vanilla CoVoGAN Varying $z^{s}$ Varying $z^{c}$
MCC (%) $\uparrow$ 33.78 36.89 31.76
SAP (%) $\uparrow$ 8.48 8.71 10.39
Modularity (%) $\uparrow$ 17.37 14.75 11.29
@@ F.7 Experiments on Independent Utility of $z^{c}$ and $z^{s}$
Figure 8: Generate same motion of different identities on Taichi Dataset.
Figure 9: Controllable text-to-video generation results.
Appendix G Impact Statements